OpenStudy (anonymous):

given y>0 and dy/dx=3x^2+4x/y. if the point (1,sqrt{10}) is on the graph relating x and y, then what is y when x=0?

OpenStudy (turingtest):

no chance that's supposed to be dy/dx=(3x^2+4x)/y is it ?

myininaya (myininaya):

it is

myininaya (myininaya):

he didn't like my answer maybe you can try turing

OpenStudy (turingtest):

Ah, I see... You should trust myininaya spikey

OpenStudy (anonymous):

lol i didnt understand it :(

myininaya (myininaya):

but no he might understand you better

OpenStudy (anonymous):

dy/dx=(3x^2+4x)/y yes thats correct

myininaya (myininaya):

this is what i did turing if you want to look

OpenStudy (anonymous):

yep thats the other one :) do you wanna work off here or of the other one.

OpenStudy (turingtest):

ok... this is a very basic separation of variables problem, in which we can treat dy/dx like a regular fraction... dy/dx=(3x^2+4x)/y ydy=3x^2+4xdx now integrate both sides... but according to the other post you don't seem to know how integrate yet :/

OpenStudy (turingtest):

you don't know how to integrate that* correct?

OpenStudy (anonymous):

hmm im not sure. if you mean using anti derivaties than sort of

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\] but i havent used this yet

OpenStudy (turingtest):

yes exactly antiderivatives

myininaya (myininaya):

\[\int\limits_{}^{}x^n dx=\frac{x^{n+1}}{n+1}+C, n \neq -1\]

myininaya (myininaya):

or you could say the antiderivative of x^n = that

OpenStudy (anonymous):

ahh yes that sort of makes sense.

OpenStudy (anonymous):

i understand that c is a constant.

OpenStudy (turingtest):

check that the derivative of that gives x^n

myininaya (myininaya):

right

OpenStudy (anonymous):

the derivative of ydy=(3x^2+4x)dx?

OpenStudy (turingtest):

antiderivative of that

myininaya (myininaya):

no we want to integrate both sides or we can use the term antiderivative

myininaya (myininaya):

like turing said

OpenStudy (anonymous):

ahh okay.

OpenStudy (turingtest):

so what is the antiderivative of ydy according to the formula myin posted?

OpenStudy (anonymous):

x^3

OpenStudy (anonymous):

for 3x^2

OpenStudy (anonymous):

i dont know for 4x

OpenStudy (turingtest):

yes but that is not ydy... like I said, I can't use latex, myin will have to show you

myininaya (myininaya):

\[\int\limits_{}^{}y dy =\int\limits_{}^{}y^1 dy=\frac{y^{1+1}}{1+1}+c_1\]

myininaya (myininaya):

see n was 1 here

OpenStudy (anonymous):

okay im struggling alot right now because i havent had enough practice with this.

myininaya (myininaya):

its cool

myininaya (myininaya):

struggling happens but the main is don't give up

OpenStudy (anonymous):

okay so the rule is for 3x^2 x^n+1

OpenStudy (anonymous):

what formula for the 4x?

OpenStudy (anonymous):

the anti derivative of dx is x?

OpenStudy (turingtest):

4x=4x^(1) so we can use the same formula integral of (x^n)dx=x^(n+1)/(n+1)

myininaya (myininaya):

\[\int\limits_{}^{}3x^2 dx=3 \int\limits_{}^{}x^2 dx=3 \cdot \frac{x^{2+1}}{2+1}+c_2\]

myininaya (myininaya):

\[\int\limits_{}^{}4x dx=4 \int\limits_{}^{}x dx=4\int\limits_{}^{}x^1 dx=4 \cdot \frac{x^{1+1}}{1+1}+c_3\]

myininaya (myininaya):

see i'm using that same formula every time

OpenStudy (anonymous):

oh okay so i have to do it for each.

myininaya (myininaya):

right

myininaya (myininaya):

\[\int\limits_{}^{}(f(x)+g(x))dx=\int\limits_{}^{}f(x) dx+\int\limits_{}^{}g(x)dx\]

myininaya (myininaya):

so we have \[\frac{y^2}{2}+c_1=3 \cdot \frac{x^{2+1}}{2+1}+c_2+4 \cdot \frac{x^{1+1}}{1+1}+c_3\]

OpenStudy (anonymous):

so for the formula ydy you used the formula.

OpenStudy (anonymous):

okay i understand how u got that equation.

myininaya (myininaya):

or you could write instead \[\frac{y^2}{2}=3 \cdot \frac{x^{2+1}}{2+1}+4 \cdot \frac{x^{1+1}}{1+1}+C\] since the sum of some constants is still a constant

OpenStudy (anonymous):

yes

myininaya (myininaya):

ok great! lets make this prettier

myininaya (myininaya):

\[\frac{y^2}{2}=x^3+2 x^2+C\]

myininaya (myininaya):

is that okay?

OpenStudy (anonymous):

yes my simplification came out the same

myininaya (myininaya):

ok we can also multiply two on both sides

myininaya (myininaya):

\[y^2=2x^3+4x^2+C\]

myininaya (myininaya):

2C is still constant so I left it as C

OpenStudy (anonymous):

yes

myininaya (myininaya):

you can write 2C if you feel more comfortable with that

OpenStudy (anonymous):

and now square both sides?

myininaya (myininaya):

square root of both sides

myininaya (myininaya):

we don't need to keep the negative value since your directions say y>0

OpenStudy (anonymous):

where do we have a negative value?

myininaya (myininaya):

take square root of both sides you get plus or minus

OpenStudy (anonymous):

thats right!

myininaya (myininaya):

we only need the plus since y>0

myininaya (myininaya):

so what I'm saying is that we have \[y=\sqrt{2x^3+4x^2+C}\]

OpenStudy (anonymous):

yes.

myininaya (myininaya):

you were given a point on this curve

myininaya (myininaya):

\[ (1,\sqrt{10} )\]

myininaya (myininaya):

x=1 and y=sqrt(10)

OpenStudy (anonymous):

yes i plug it in for x and y?

myininaya (myininaya):

so we can use this to find C

myininaya (myininaya):

\[\sqrt{10}=\sqrt{2 (1)^3+4(1)^2+C}\] => \[10=2(1)^3+4(1)^2+C\]

myininaya (myininaya):

\[10=2+4+C\]

myininaya (myininaya):

=>C=4

myininaya (myininaya):

so we have \[y=\sqrt{2x^3+4x^2+4} \]

myininaya (myininaya):

you wanted to know y when x=0, right?

OpenStudy (anonymous):

yes

myininaya (myininaya):

so how do you think we do that?

myininaya (myininaya):

i brb i think it will be pretty easy for turing to help without latex on this last part that you have to do

OpenStudy (anonymous):

okay turing i get now :D

OpenStudy (anonymous):

but i need help finishing

myininaya (myininaya):

ok i'm back

myininaya (myininaya):

i had to get my glasses

myininaya (myininaya):

so i can be fully nerd

OpenStudy (anonymous):

lol

myininaya (myininaya):

so we have \[y=\sqrt{2x^3+4x^2+4}\]

OpenStudy (anonymous):

yes

myininaya (myininaya):

it says what is y if x=0

myininaya (myininaya):

\[y=\sqrt{2(0)^3+4(0)^2+4}\]

myininaya (myininaya):

i replaced x with 0 so I can see what y is when x is 0

myininaya (myininaya):

\[y=\sqrt{0+0+4}=\sqrt{4}=2\]

myininaya (myininaya):

This says when x is 0, y is 2

OpenStudy (anonymous):

got it! :D thank you!!!! sorry i didnt really try hard enough before you were a great help!

myininaya (myininaya):

It is okay. I didn't think any offense to anything you did. Sometimes you may get someone who can explain it better. I know that I'm probably not the best explaining some things.

myininaya (myininaya):

Or you know like I way you prefer.